OpenStudy (anonymous):

Which of the following points are NOT on the graph of y = -4x + 4? A.1, 0) B. (2, 4) C. (0, 4) D. (-1, 8)

OpenStudy (cwrw238):

just plug in each value of x and y into the function and see which fits

OpenStudy (cwrw238):

eg for A:- x = 1 and y = 0

OpenStudy (anonymous):

just want to check an answer

OpenStudy (anonymous):

i beleive it is C

OpenStudy (cwrw238):

C 4 = -4(0) + 4 = 4 not C

OpenStudy (anonymous):

what are you getting

OpenStudy (cwrw238):

try b

OpenStudy (cwrw238):

4 = -4(2) + 4 ? are they equal?

OpenStudy (anonymous):

no

OpenStudy (cwrw238):

so thats it

OpenStudy (anonymous):

what letter is that

OpenStudy (cwrw238):

B

OpenStudy (anonymous):

ok thx

OpenStudy (cwrw238):

ok